# Borel 1 type mappings and the respective equi-families

### v2

@@ 2.1. Relations between various families of functions
$n(x):=\min\{n:x\in X_{n}\}\textup{\ and\ }V_{x}:=X\setminus\bigcup_{n<n(x)}X_{n}\setminus\bigcup\{F\in\mathcal{A}_{n(x)}\colon x\notin F\}.$
@@
We show that (GLP) $\Rightarrow$ (Borel 1). Let $f\colon X\to Y$ have the generalized Lebesgue property. For every $n\in\mathbb{N}$ we choose a family $\mathcal{A}$ of subsets of $X$ such that $\mathcal{A}_{n}=\bigcup_{k\in\mathbb{N}}\mathcal{A}_{n,k}$, where each $\mathcal{A}_{n,k}$ is a discrete family consisting of closed sets, $\mathcal{A}_{n}$ is a covering of $X$ and ${\rm diam}f(A)<1/n$ for each $A\in\mathcal{A}_{n}$. Denote $\mathcal{B}:=\bigcup_{n,k\in\mathbb{N}}\mathcal{A}_{n,k}$. Notice that for every subfamily $\mathcal{B}^{\prime}\subseteq\mathcal{B}$ the union $\bigcup\mathcal{B}^{\prime}$ of all members in $\mathcal{B}^{\prime}$ is an $F_{\sigma}$-set, since every $\mathcal{A}_{k,n}$ is a discrete family of closed sets. It is easy to see that
@@
Finally, we show that (GLP) $\Rightarrow$ (Cliq) if $X$ is a Baire space (see also [17, proof of Theorem 3.8]). Let $f\colon X\to Y$ have the generalized Lebesgue property, $\varepsilon>0$ and $\emptyset\neq U\subseteq X$ be an open set. Similarly as in the proof of “(GLP) $\Rightarrow$ (WS)” there exists a closed cover $\mathcal{A}_{\varepsilon}$ of $X$ such that ${\rm diam}f(A)\leq\varepsilon$ for all $A\in\mathcal{A}$, and $\mathcal{A}=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}$, where $\mathcal{A}_{n}$ is discrete for each $n\in\mathbb{N}$. Note that for all $n\in\mathbb{N}$ and $x\in\bigcup\mathcal{A}_{n}$ there exists a unique $F^{n}_{x}\in\mathcal{A}_{n}$ such that $x\in F^{n}_{x}$. Let $X_{n}:=\bigcup\mathcal{A}_{n}$ for each $n\in\mathbb{N}$. Note that $(X_{n})_{n\in\mathbb{N}}$ is a cover of $X$ consisting of closed sets. Since $U=\bigcup_{n\in\mathbb{N}}(X_{n}\cap U)$ and $U$ is a Baire space, there is a nonempty open set $V\subseteq U$ and $n_{0}\in\mathbb{N}$ such that $V\subseteq X_{n_{0}}\cap U$. Fix $x\in V$. Since $\mathcal{A}_{n_{0}}$ is discrete, there exists a nonempty open set $V^{\prime}$ such that $V^{\prime}\cap X_{n_{0}}=V^{\prime}\cap F^{x}_{n_{0}}$. Then ${\rm diam}f(V\cap V^{\prime})\leq\varepsilon$.
@@ 3.1. Equi-families and the orbit function
  • $\mathcal{F}$ is equi-fragmentable, if for every $\varepsilon>0$ and every closed set $\emptyset\neq F\subseteq X$ there exists an open set $U$ such that $U\cap F\neq\emptyset$ and ${\rm diam}f(U\cap F)<\varepsilon$ and for all $f\in\mathcal{F}$;
@@ 3.2. Closures in the topology of pointwise convergence
$(x,y)\in V_{y}\times V_{x}\textup{\ implies\ }d(f(x),f(y))\leq\varepsilon.$
@@ 3.3. Applications to functions defined on $X\times Y$
for all $A\in\mathcal{A}$ and for all $x\in X$. Let $\mathcal{A}=\bigcup\limits_{n}\mathcal{A}_{n}$, where each family $\mathcal{A}_{n}$ is discrete in $Y$. We put
$\mathcal{B}_{n}=(A\setminus\bigcup\limits_{k<n}\bigcup\mathcal{A}_{k}:A\in\mathcal{A}_{n}),\,\,n\in\mathbb{N},$
$B_{n}=\bigcup\{B:B\in\mathcal{B}_{n}\}$ and  $\mathcal{B}=\bigcup\limits_{n=1}^{\infty}\mathcal{B}_{n}.$
@@
Take an arbitrary $y_{B,n}\in B$ for every $B\in\mathcal{B}_{n}$, $n\in\mathbb{N}$. For each $y\in Y$ we put $n(y)=\min\{n\in\mathbb{N}:y\in B_{n}\}$. Now let $(x,y)\in X\times Y$. Since the family $\mathcal{B}_{n(y)}$ is disjoint, there exists a unique set $B(y)\in\mathcal{B}_{n(y)}$ such that $y\in B(y)$. We define

### v3

@@ 2.1. Relations between various families of functions
$n(x):=\min\{n:x\in X_{n}\}\textup{\ and\ }V_{x}:=X\setminus\bigcup_{n<n(x)}\left(X_{n}\setminus\bigcup\{F\in\mathcal{A}_{n(x)}\colon x\notin F\}\right).$
@@
We show that (GLP) $\Rightarrow$ (Borel 1). Let $f\colon X\to Y$ have the generalized Lebesgue property. For every $n\in\mathbb{N}$ we choose a family $\mathcal{A}_{n}$ of subsets of $X$ such that $\mathcal{A}_{n}=\bigcup_{k\in\mathbb{N}}\mathcal{A}_{n,k}$, where each $\mathcal{A}_{n,k}$ is a discrete family consisting of closed sets, $\mathcal{A}_{n}$ is a covering of $X$ and ${\rm diam}f(A)<1/n$ for each $A\in\mathcal{A}_{n}$. Denote $\mathcal{B}:=\bigcup_{n,k\in\mathbb{N}}\mathcal{A}_{n,k}$. Notice that for every subfamily $\mathcal{B}^{\prime}\subseteq\mathcal{B}$ the union $\bigcup\mathcal{B}^{\prime}$ of all members in $\mathcal{B}^{\prime}$ is an $F_{\sigma}$-set, since every $\mathcal{A}_{k,n}$ is a discrete family of closed sets. It is easy to see that
@@
Finally, we show that (GLP) $\Rightarrow$ (Cliq) if $X$ is a Baire space (see also [17, proof of Theorem 3.8]). Let $f\colon X\to Y$ have the generalized Lebesgue property, $\varepsilon>0$ and $\emptyset\neq U\subseteq X$ be an open set. Similarly as in the proof of “(GLP) $\Rightarrow$ (WS)” there exists a closed cover $\mathcal{A}$ of $X$ such that ${\rm diam}f(A)\leq\varepsilon$ for all $A\in\mathcal{A}$, and $\mathcal{A}=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}$, where $\mathcal{A}_{n}$ is discrete for each $n\in\mathbb{N}$. Note that for all $n\in\mathbb{N}$ and $x\in\bigcup\mathcal{A}_{n}$ there exists a unique $F^{n}_{x}\in\mathcal{A}_{n}$ such that $x\in F^{n}_{x}$. Let $X_{n}:=\bigcup\mathcal{A}_{n}$ for each $n\in\mathbb{N}$. Note that $(X_{n})_{n\in\mathbb{N}}$ is a cover of $X$ consisting of closed sets. Since $U=\bigcup_{n\in\mathbb{N}}(X_{n}\cap U)$ and $U$ is a Baire space, there is a nonempty open set $V\subseteq U$ and $n_{0}\in\mathbb{N}$ such that $V\subseteq X_{n_{0}}\cap U$. Fix $x\in V$. Since $\mathcal{A}_{n_{0}}$ is discrete, there exists a nonempty open set $V^{\prime}$ such that $V^{\prime}\cap X_{n_{0}}=V^{\prime}\cap F^{x}_{n_{0}}$. Then ${\rm diam}f(V\cap V^{\prime})\leq\varepsilon$.
@@ 3.1. Equi-families and the orbit function
  • $\mathcal{F}$ is equi-fragmentable, if for every $\varepsilon>0$ and every closed set $\emptyset\neq F\subseteq X$ there exists an open set $U$ such that $U\cap F\neq\emptyset$ and ${\rm diam}f(U\cap F)<\varepsilon$ for all $f\in\mathcal{F}$;
@@ 3.2. Closures in the topology of pointwise convergence
$(x,y)\in V_{y}\times V_{x}\textup{\ implies\ }d(g(x),g(y))\leq\varepsilon.$
@@ 3.3. Applications to functions defined on $X\times Y$
for all $A\in\mathcal{A}$ and $x\in X$. Let $\mathcal{A}=\bigcup\limits_{n}\mathcal{A}_{n}$, where each family $\mathcal{A}_{n}$ is discrete in $Y$. For every $n\in\mathbb{N}$ we put
$\mathcal{B}_{n}=\{A\setminus\bigcup\limits_{k<n}\bigcup\mathcal{A}_{k}:A\in\mathcal{A}_{n}\}$   and   $B_{n}=\bigcup\{B:B\in\mathcal{B}_{n}\}$.
Notice that each family $\mathcal{B}_{n}$ is discrete (in particular, disjoint), since $\mathcal{B}_{n}\prec\mathcal{A}_{n}$. Let
$\mathcal{B}=\bigcup\limits_{n=1}^{\infty}\mathcal{B}_{n}.$
@@
Take an arbitrary $y_{B,n}\in B$ for every nonempty $B\in\mathcal{B}_{n}$, $n\in\mathbb{N}$. For each $y\in Y$ we put $n(y)=\min\{n\in\mathbb{N}:y\in B_{n}\}$. Now let $(x,y)\in X\times Y$. Since the family $\mathcal{B}_{n(y)}$ is disjoint, there exists a unique set $B(y)\in\mathcal{B}_{n(y)}$ such that $y\in B(y)$. We define